OpenStudy (anonymous):

What is the electric force between a +2µC point charge and a -2µC point charge if they are separated by a distance of 5.0 cm? Show your work. (µC=1.0x10^-6C) Please HELP. Will give medal.

OpenStudy (isaiah.feynman):

The electric force between two charges is given by \[F = K\frac{ q_{1}{q_{2}} }{ d^{2} }\]

OpenStudy (anonymous):

so how will I plug that in?

OpenStudy (isaiah.feynman):

Q1 and Q2 are the charges, d is the distance between the charges. K is a special constant, it given as \[K = \frac{ 1 }{ 4\pi \epsilon_{o} } \]So, I can write the first equation as \[F = \left( \frac{ 1 }{ 4\pi \epsilon_{o} } \right)\left( \frac{ q_{1}q_{2} }{ d^{2} } \right)\]

OpenStudy (anonymous):

I figured out the answer thank you for your help!!